# Solution of some tiling open problems of Propp, Lai, and some related results

### v2

@@ 3. Solution of Four Open Problems of Propp
In this section we solve four open problems posed by Propp [29] on the enumeration of trimer coverings of regions called benzels. To state the problems, we first define the benzel regions. For any positive integers $a$ and $b$ such that $2\leq a\leq 2b$ and $2\leq b\leq 2a$, we defined the $(a,b)$-benzel as follows. Consider a hexagonal grid on the plane consisting of regular hexagons of side-length $1$, each having two horizontal sides (see Figure 7; this can be thought of as a tessellation of the plane by unit hexagons). We call these unit hexagons cells. Regard the plane as the complex plane so that the origin is the left vertex of one of the cells. Let $H$ be the hexagon whose vertices are $\omega^{j}(a\omega+b)$ and $-\omega^{j}(a+b\omega)$ for $j\in\{0,1,2\}$, where $\omega=e^{2\pi i/3}$. The $(a,b)$-benzel is defined to be the union of all the cells that are contained in the hexagon $H$; Figure 7 shows the $(7,8)$-benzel.
@@
One example is: “How many trimer covers does the $(a,b)$-benzel have, in which all the three types of bones, but not both types of stones are allowed?” (Problem 3 in [29]). See [13][14][20] for recent progress on some of these problems. The open problems that we solve in this section are about the enumeration of trimer covers of benzels that can use both kinds of stones, the rising bone and the falling bone (in other words, the only restriction is that they cannot use any vertical bone). In this paper, we call such tilings vertical-bone-free tilings. Based on numerical data, Propp conjectured that the number of such trimer covers of the $(3n,3n)$-benzel (Problem 8), the $(3n+1,3n+1)$-benzel (Problem 9), the $(3n+1,3n+2)$-benzel (Problem 10), and the $(3n-1,3n)$-benzel (Problem 11) satisfy explicit recurrence relations. Proving such statements is in principle very hard, as, in contrast to the situation of dimer coverings, there are not many enumeration techniques available for trimer covers.
@@
In [13] Defant et al. proved that the trimer enumeration questions in Problems 8–11 on Propp’s list can be restated in terms of domino tilings of truncated Aztec diamonds as follows101010 For a lattice region $R$ on the square lattice, $\operatorname{M}(R)$ denotes the number of its domino tilings. (see Conjectures 5.1–5.3 and Question 5.4 in [13]).
@@
((((a)))). The number of vertical-bone-free tilings of the $(3n,3n)$-benzel is $\operatorname{M}(TAD_{2n-1})$.
((((b)))). The number of vertical-bone-free tilings of the $(3n+1,3n+1)$-benzel is $\operatorname{M}(TAD_{2n})$.
((((c)))). The number of vertical-bone-free tilings of the $(3n-1,3n)$-benzel is $\operatorname{M}(TAD^{\prime}_{2n-1})$.
((((d)))). The number of vertical-bone-free tilings of the $(3n+1,3n+2)$-benzel is $\operatorname{M}(TAD^{\prime}_{2n})$.
@@
For positive integers $n$, the numbers of domino tilings of the regions $TAD_{2n-1}$, $TAD_{2n}$, $TAD^{\prime}_{2n-1}$, and $TAD^{\prime}_{2n}$ are given by the following product formulas111111 Throughout this paper, an empty product is defined to be $1$ and $(2m-1)!!\coloneqq(2m-1)(2m-3)\cdots 1$ for postive integers $m$.:
@@
((((a)))). The number of vertical-bone-free tilings of $(3n,3n)$-benzel is given by (3.1).
((((b)))). The number of vertical-bone-free tilings of $(3n+1,3n+1)$-benzel is given by (3.2).
((((c)))). The number of vertical-bone-free tilings of $(3n-1,3n)$-benzel is given by (3.3).
((((d)))). The number of vertical-bone-free tilings of $(3n+1,3n+2)$-benzel is given by (3.4).
@@
One of these families consists of the Aztec triangles introduced by Di Francesco [15], and the remaining three are variants of it. All of them can be regarded as hybrids between Aztec diamonds and grid squares. Namely, let $S_{2n}$ be the square grid of side-length $2n$. Cut it in two congruent parts by a zig-zag lattice path of step length 2, as shown in Figure 12, and glue together the top half with the top half of $AD_{n-1}$. If this is done in a right-justified manner, one obtains Di Francesco’s Aztec triangle ${\mathcal{T}}_{n}$ (see the picture in the center in Figure 12); the left-justified fashion results in the Aztec triangle ${\mathcal{T}}^{\prime}_{n}$ (see the picture on the right in Figure 12).
@@
These four families of regions turn out to be special cases of regions considered by Corteel et al. in [12]. In that paper, the authors consider a family of regions parametrized by positive integers $l$ and $k$. The region ${\mathcal{T}}_{n}$ represents the special case $l=2n$ and $k=n$. Furthermore, ${\mathcal{T}}^{\prime}_{n}$ is the special case $l=2n+1$ and $k=n-1$, ${\mathcal{T}}^{\prime\prime}_{n}$ the special case $l=2n+1$ and $k=n$, and ${\mathcal{T}}^{\prime\prime\prime}_{n}$ is obtained by setting $l=2n$ and $k=n-1$.
@@
$\operatorname{M}({\mathcal{T}}^{\prime\prime\prime}_{n})=2^{n(n-1)/2}\cdot\frac{(2n-1)!!}{n!}\cdot\prod_{i=0}^{n-1}\frac{(4i)!}{(n+2i)!}.$ (3.7)
@@
Remark $2$. The original conjectures of Propp did not state explicit formulas for the number of vertical-bone-free tilings of benzels. Instead, the conjectures stated that those numbers satisfy certain recurrence relations. More precisely, via the compression bijection of [13], Propp’s conjectures (open Problems 8–11 in [29]) are equivalent to the following equalities ((3.12) and (3.15) were given in [29] and the other two in [13]):
@@ 4.1. Six families of hexagonal regions with collinear holes
Let $\ell$ be the vertical symmetry axis of $H(a,b,k)$, and label the slots along $\ell$ choosing the reference line $L$ to be the base of $H(a,b,k)$ (see the picture on the left in Figure 16 for an example). Choose an arbitrary subset of the labeled slots along $\ell$ (avoiding the possible partial slot at the top), and for each chosen slot remove the up-pointing lattice triangle of side-length two containing that slot. If the resulting region has forced lozenges, remove them121212 By this, if a run of $k$ triangular holes of side-length two are contiguous, this gets replaced by a single triangular hole of side-length $2k$. In addition, it follows that the top triangular hole in $H_{\mathbf{l}}(a,b,k)$ does not touch the top side of the boundary.. Let ${\mathbf{l}}$ be the list of the labels of slots not covered by removed triangles in the resulting region. Then we denote the resulting region by $H_{\mathbf{l}}(a,b,k)$; an example is shown on the left in Figure 16.
@@
The last pair of families of regions is very similar to the previous one; the only difference is that the triangular hole of odd side-length is chosen to point downward131313 Note that such a region cannot be obtained by rotating an $H_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)$ by $180^{\circ}$, because we are assuming that $k$, which is the difference between the number of up-pointing and down-pointing unit triangles in $H(a,b,k)$, is non-negative.. We denote the resulting regions by $\bar{H}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)$ and $\bar{H}^{\prime}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)$, respectively (these are illustrated in Figure 18).
@@ 4.2. Two families of dented regions
Given non-empty lists of strictly increasing positive integers ${\mathbf{l}}=(l_{1},\dotsc,l_{m})$ and ${\mathbf{q}}=(q_{1},\dotsc,q_{n})$ and a non-negative integer $x$, the region $R_{{\mathbf{l}},{\mathbf{q}}}(x)$ is defined as indicated in Figure 20. The lists ${\mathbf{l}}$ and ${\mathbf{q}}$ specify which bumps on the zigzag lines are kept. The kept bumps on $P_{u}$ are joined together by creating up-pointing dents, and those on $P_{d}$ by down-pointing dents; these two portions are then joined together via a horizontal ray left of $O$, as shown; $x$ is the length of the base. Note that this information, together with the fact that we want a region which can be tiled by lozenges, determines the rest of the boundary of $R_{{\mathbf{l}},{\mathbf{q}}}(x)$141414 Indeed, if a lozenge tiling exists, due to a family of paths of lozenges that it determines, the length of the southwest side must be equal to the number of unit segments facing northeast on the right boundary..
In case one of the lists ${\mathbf{l}}$ or ${\mathbf{q}}$ is empty, the definition is slightly different, as shown in Figure 20; note in particular that if ${\mathbf{l}}=\emptyset$, $x$ is not the length of the base, but one unit less than the distance from the left end of the base to the reference point $O$. Define $R_{\emptyset,\emptyset}(x)=\emptyset$ for all $x$.
@@
Our second family of regions, denoted $\bar{R}_{{\mathbf{l}},{\mathbf{q}}}(x)$, is defined almost identically (see Figures 22 and 22). The only difference is that, when connecting the upper and lower boundary portions determined by the selected bumps, instead of the horizontal ray starting from $O$, we use the one starting from the lattice point $\bar{O}$, one step southwest of $O$ (and, in case ${\mathbf{l}}$ is empty, when defining the bottom side of $\bar{R}_{\emptyset,{\mathbf{q}}}(x)$, we move left until we are $x$ units away from $\bar{O}$). Again, we define $\bar{R}_{\emptyset,\emptyset}(x)=\emptyset$ for all $x$.
@@ 4.3. Two families of polynomials and their connection to the dented regions
×$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \times\prod_{i=1}^{n}\frac{(x+i)_{m}}{(x+i+1/2)_{m}}\prod_{i=1}^{m}(2x+n+i+2)_{n+i-1}$
@@
×$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \times\prod_{i=1}^{m}\frac{(x+i)_{n}}{(x+i+1/2)_{n}}\prod_{i=1}^{n}(2x+m+i+1)_{m+i}.$
@@
We are now ready to define our polynomials $P_{\mathbf{l},\mathbf{q}}(x)$ and $\bar{P}_{\mathbf{l},\mathbf{q}}(x)$. For lists of strictly increasing positive integers ${\mathbf{l}}=(l_{1},\ldots,l_{m})$ and ${\mathbf{q}}=(q_{1},\ldots,q_{n})$, $P_{\mathbf{l},\mathbf{q}}(x)$ and $\bar{P}_{\mathbf{l},\mathbf{q}}(x)$ are given by  [^15] These are equations (5.1) and (5.2) in [5, Part B].161616 For $m=0$, one needs to take $l_{m}=0$ in equations (4.1) and (4.2); also, when $m$ or $n$ is zero at the upper limit of an outside product, that double product is defined to be 1.
@@
Let $H$ be one of the regions $H_{\mathbf{l}}(a,b,k)$, $H_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)$ or $\bar{H}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)$. Then we have171717 Here $H^{+}$ and $H^{-}$ are the regions obtained from the symmetric region $H$ by the procedure described before the statement of Theorem 2.1.
@@
and all of the regions $H^{+}$ and $H^{-}$ are $R$- or $\bar{R}$-regions, as follows181818 Given two regions $R$ and $Q$, $[R,Q]$ is the (ordered) list consisting of these regions.:
((((a)))). For $k$ even, we have
$[H_{\mathbf{l}}(a,b,k)^{+},H_{\mathbf{l}}(a,b,k)^{-}]=\begin{cases}[\bar{R}_{{\mathbf{l}}-1,\emptyset}(a/2),R_{\emptyset,{\mathbf{l}}}((a+k-2)/2)],&\text{$a$ even, $l_{1}=1$,}\\ [{R}_{{\mathbf{l}}-1,\emptyset}(a/2),R_{\emptyset,{\mathbf{l}}}((a+k-2)/2)],&\text{$a$ even, $l_{1}>1$,}\\ [\bar{R}_{{\mathbf{l}},\emptyset}((a-1)/2),R_{\emptyset,{\mathbf{l}}^{(m)}}((a+k-1)/2)],&\text{$a$ odd,}\end{cases}$ (4.6)
where ${\mathbf{l}}-1$ denotes the list obtained from ${\mathbf{l}}$ by decrementing each of its elements by one unit ((((discarding the first element if $l_{1}=1$)))).
((((b)))). For $k$ odd, we have
$[H_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)^{+},H_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)^{-}]=\begin{dcases}[{R}_{{\mathbf{q}},{\mathbf{l}}^{(m)}}(a/2),\bar{R}_{{\mathbf{l}},{\mathbf{q}}}((a+k-1)/2)],&\text{$a$ even,}\\ [{R}_{{\mathbf{q}},{\mathbf{l}}}((a-1)/2),\bar{R}_{{\mathbf{l}},{\mathbf{q}}^{(n)}}((a+k)/2)],&\text{$a$ odd,}\end{dcases}$ (4.7)
@@
$[\bar{H}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)^{+},\bar{H}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)^{-}]=\begin{dcases}[\bar{R}_{{\mathbf{q}},{\mathbf{l}}^{(m)}}(a/2),{R}_{{\mathbf{l}},{\mathbf{q}}}((a+k-1)/2)],&\text{$a$ even,}\\ [\bar{R}_{{\mathbf{q}},{\mathbf{l}}}((a-1)/2),{R}_{{\mathbf{l}},{\mathbf{q}}^{(n)}}((a+k)/2)],&\text{$a$ odd,}\end{dcases}$ (4.8)
@@
Let $H$ be one of the regions $H^{\prime}_{\mathbf{l}}(a,b,k)$, $H^{\prime}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)$ or $\bar{H}^{\prime}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)$, and let $\ell^{\prime}$ be the axis with respect to which the holes in $H$ are symmetric. Then we have191919 Here ${H}^{+}$ and ${H}^{-}$ are the “half-regions” obtained from the nearly symmetric hexagon with holes $H$ by the procedure described before the statement of Theorem 2.2 (see the middle picture in Figure 23 for an example).
@@
((((a)))). For $k$ even, we have
$[H^{\prime}_{\mathbf{l}}(a,b,k)^{+},H^{\prime}_{\mathbf{l}}(a,b,k)^{-}]=\begin{cases}[\bar{R}_{{\mathbf{l}}-1,\emptyset}((a-1)/2),R_{\emptyset,{\mathbf{l}}}((a+k-1)/2)],&\text{$a$ odd, $l_{1}=1$,}\\ [R_{{\mathbf{l}}-1,\emptyset}((a-1)/2),R_{\emptyset,{\mathbf{l}}}((a+k-1)/2)],&\text{$a$ odd, $l_{1}>1$,}\\ [\bar{R}_{{\mathbf{l}},\emptyset}((a-2)/2),R_{\emptyset,{\mathbf{l}}^{(m)}}((a+k)/2)],&\text{$a$ even.}\end{cases}$ (4.10)
((((b)))). For $k$ odd, we have
$[H^{\prime}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)^{+},H^{\prime}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)^{-}]=\begin{dcases}[{R}_{{\mathbf{q}},{\mathbf{l}}^{(m)}}((a-1)/2),\bar{R}_{{\mathbf{l}},{\mathbf{q}}}((a+k)/2)],&\text{$a$ odd,}\\ [{R}_{{\mathbf{q}},{\mathbf{l}}}((a-2)/2),\bar{R}_{{\mathbf{l}},{\mathbf{q}}^{(n)}}((a+k+1)/2)],&\text{$a$ even,}\end{dcases}$ (4.11)
@@
$[\bar{H}^{\prime}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)^{+},\bar{H}^{\prime}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)^{-}]=\begin{dcases}[\bar{R}_{{\mathbf{q}},{\mathbf{l}}^{(m)}}((a-1)/2),{R}_{{\mathbf{l}},{\mathbf{q}}}((a+k)/2)],&\text{$a$ odd,}\\ [\bar{R}_{{\mathbf{q}},{\mathbf{l}}}((a-2)/2),{R}_{{\mathbf{l}},{\mathbf{q}}^{(n)}}((a+k+1)/2)],&\text{$a$ even.}\end{dcases}$ (4.12)
@@
Let $H$ be one of the regions $H^{\prime}_{\mathbf{l}}(a,b,k)$, $H^{\prime}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)$ or $\bar{H}^{\prime}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)$, and let $\ell^{\prime}$ be the axis with respect to which the holes in $H$ are symmetric. Then we have202020 Here $\widehat{H}^{+}$ and $\widehat{H}^{-}$ are the “half-regions” obtained from the nearly symmetric hexagon with holes $H$ by the procedure described before the statement of Theorem 2.2 (see the middle picture in Figure 24 for an example).

### v3

@@ 3. Solution of Four Open Problems of Propp
In this section we solve four open problems posed by Propp [30] on the enumeration of trimer coverings of regions called benzels. To state the problems, we first define the benzel regions. For any positive integers $a$ and $b$ such that $2\leq a\leq 2b$ and $2\leq b\leq 2a$, we defined the $(a,b)$-benzel as follows. Consider a hexagonal grid on the plane consisting of regular hexagons of side-length $1$, each having two horizontal sides (see Figure 9; this can be thought of as a tessellation of the plane by unit hexagons). We call these unit hexagons cells. Regard the plane as the complex plane so that the origin is the left vertex of one of the cells. Let $H$ be the hexagon whose vertices are $\omega^{j}(a\omega+b)$ and $\omega^{j}(a+b\omega)$ for $j\in\{0,1,2\}$, where $\omega=e^{2\pi i/3}$. The $(a,b)$-benzel is defined to be the union of all the cells that are contained in the hexagon $H$; Figure 9 shows the $(7,8)$-benzel.
@@
One example is: “How many trimer covers does the $(a,b)$-benzel have, in which all the three types of bones, but not both types of stones are allowed?” (Problem 3 in [30]). See [13][14][20] for recent progress on some of these problems. The open problems that we solve in this section are about the enumeration of trimer covers of benzels that can use both kinds of stones, the rising bone and the falling bone (in other words, the only restriction is that they cannot use any vertical bone). In this paper, we call such tilings vertical-bone-free tilings. Based on numerical data, Propp conjectured that the number of such trimer covers of the $(3n,3n)$-benzel (Problem 8), the $(3n+1,3n+1)$-benzel (Problem 9), the $(3n+1,3n+2)$-benzel (Problem 10), and the (3n1,3n)(3n-1,3n)-benzel (Problem 11) satisfy explicit recurrence relations. Proving such statements is in principle very hard, as, in contrast to the situation of dimer coverings, there are not many enumeration techniques available for trimer covers.
@@
In [13] Defant et al. proved that the trimer enumeration questions in Problems 8–11 on Propp’s list can be restated in terms of domino tilings of truncated Aztec diamonds as follows111111 For a lattice region $R$ on the square lattice, $\operatorname{M}(R)$ denotes the number of its domino tilings. (see Conjectures 5.1–5.3 and Question 5.4 in [13]).
@@
((a)). The number of vertical-bone-free tilings of the $(3n,3n)$-benzel is $\operatorname{M}(TAD_{2n-1})$.
((b)). The number of vertical-bone-free tilings of the $(3n+1,3n+1)$-benzel is $\operatorname{M}(TAD_{2n})$.
((c)). The number of vertical-bone-free tilings of the (3n1,3n)(3n-1,3n)-benzel is $\operatorname{M}(TAD^{\prime}_{2n-1})$.
((d)). The number of vertical-bone-free tilings of the $(3n+1,3n+2)$-benzel is $\operatorname{M}(TAD^{\prime}_{2n})$.
@@
For positive integers $n$, the numbers of domino tilings of the regions $TAD_{2n-1}$, $TAD_{2n}$, $TAD^{\prime}_{2n-1}$, and $TAD^{\prime}_{2n}$ are given by the following product formulas121212 Throughout this paper, an empty product is defined to be $1$ and $(2m-1)!!\coloneqq(2m-1)(2m-3)\cdots 1$ for postive integers $m$.:
@@
((a)). The number of vertical-bone-free tilings of the $(3n,3n)$-benzel is given by (3.1).
((b)). The number of vertical-bone-free tilings of the $(3n+1,3n+1)$-benzel is given by (3.2).
((c)). The number of vertical-bone-free tilings of the (3n1,3n)(3n-1,3n)-benzel is given by (3.3).
((d)). The number of vertical-bone-free tilings of the $(3n+1,3n+2)$-benzel is given by (3.4).
@@
One of these families consists of the Aztec triangles introduced by Di Francesco [15], and the remaining three are variants of it. All of them can be regarded as hybrids between Aztec diamonds and squares. Namely, let $S_{2n}$ be the square of side-length $2n$ drawn on the square grid. Cut it in two congruent parts131313 This way of cutting up the square appears also in the second author’s paper[4] and in Pachter’s work [29]. by a zig-zag lattice path of step length 2, as shown in Figure 14, and glue together the top half with the top half of $AD_{n-1}$. If this is done in a right-justified manner, one obtains Di Francesco’s Aztec triangle ${\mathcal{T}}_{n}$ (see the picture in the center in Figure 14); the left-justified fashion results in the Aztec triangle ${\mathcal{T}}^{\prime}_{n}$ (see the picture on the right in Figure 14).
@@
These four families of regions turn out to be special cases of regions considered by Corteel et al. in [12]. In that paper, the authors consider a family of regions parametrized by positive integers $l$ and $k$. The region ${\mathcal{T}}_{n}$ represents the special case $l=2n$ and $k=n$. Furthermore, ${\mathcal{T}}^{\prime}_{n}$ is the special case $l=2n+1$ and k=n1k=n-1, ${\mathcal{T}}^{\prime\prime}_{n}$ the special case $l=2n+1$ and $k=n$, and ${\mathcal{T}}^{\prime\prime\prime}_{n}$ is obtained by setting $l=2n$ and k=n1k=n-1.
@@
$\operatorname{M}({\mathcal{T}}^{\prime\prime\prime}_{n})=2^{n(n-1)/2}\cdot\frac{(2n-1)!!}{n!}\cdot\prod_{i=1}^{n-1}\frac{(4i)!}{(n+2i)!}.$ (3.7)
@@
Remark 22. The original conjectures of Propp did not state explicit formulas for the number of vertical-bone-free tilings of benzels. Instead, the conjectures stated that those numbers satisfy certain recurrence relations. More precisely, via the compression bijection of [13], Propp’s conjectures (open Problems 8–11 in [30]) are equivalent to the following equalities ((3.12) and (3.15) were given in [30] and the other two in [13]):
@@ 4.1. Six families of hexagonal regions with collinear holes
Let $\ell$ be the vertical symmetry axis of $H(a,b,k)$, and label the slots along $\ell$ choosing the reference line $L$ to be the base of $H(a,b,k)$ (see the picture on the left in Figure 18 for an example). Choose an arbitrary subset of the labeled slots along $\ell$ (avoiding the possible partial slot at the top), and for each chosen slot remove the up-pointing lattice triangle of side-length two containing that slot. If the resulting region has forced lozenges, remove them141414 By this, if a run of $k$ triangular holes of side-length two are contiguous, this gets replaced by a single triangular hole of side-length $2k$. In addition, it follows that the top triangular hole in $H_{\mathbf{l}}(a,b,k)$ does not touch the top side of the boundary.. Let $\mathbf{l}$} be the list of the labels of slots not covered by removed triangles in the resulting region. Then we denote the resulting region by $H_{\mathbf{l}}(a,b,k)$; an example is shown on the left in Figure 18.
@@
The last pair of families of regions is very similar to the previous one; the only difference is that the triangular hole of odd side-length is chosen to point downward151515 Note that such a region cannot be obtained by rotating an $H_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)$ by $180^{\circ}$, because we are assuming that $k$, which is the difference between the number of up-pointing and down-pointing unit triangles in $H(a,b,k)$, is non-negative.. We denote the resulting regions by $\bar{H}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)$ and $\bar{H}^{\prime}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)$, respectively (these are illustrated in Figure 20).
@@ 4.2. Two families of dented regions
Given non-empty lists of strictly increasing positive integers ${\mathbf{l}}=(l_{1},\dotsc,l_{m})$ and ${\mathbf{q}}=(q_{1},\dotsc,q_{n})$ and a non-negative integer $x$, the region $R_{{\mathbf{l}},{\mathbf{q}}}(x)$ is defined as indicated in Figure 22. The lists $\mathbf{l}$} and $\mathbf{q}$} specify which bumps on the zigzag lines are kept. The kept bumps on $P_{u}$ are joined together by creating up-pointing dents, and those on $P_{d}$ by down-pointing dents; these two portions are then joined together via a horizontal ray left of $O$, as shown; $x$ is the length of the base. Note that this information, together with the fact that we want a region which can be tiled by lozenges, determines the rest of the boundary of $R_{{\mathbf{l}},{\mathbf{q}}}(x)$161616 Indeed, if a lozenge tiling exists, due to a family of paths of lozenges that it determines, the length of the southwest side must be equal to the number of unit segments facing northeast on the right boundary..
In case one of the lists $\mathbf{l}$} or $\mathbf{q}$} is empty, the definition is slightly different, as shown in Figure 22; note in particular that if $\mathbf{l}$$\emptyset$, $x$ is not the length of the base, but one unit less than the distance from the left end of the base to the reference point $O$. Define $R_{\emptyset,\emptyset}(x)=\emptyset$ for all $x$.
@@
Our second family of regions, denoted $\bar{R}_{{\mathbf{l}},{\mathbf{q}}}(x)$, is defined almost identically (see Figures 24 and 24). The only difference is that, when connecting the upper and lower boundary portions determined by the selected bumps, instead of the horizontal ray starting from $O$, we use the one starting from the lattice point $\bar{O}$, one step southwest of $O$ (and, in case $\mathbf{l}$} is empty, when defining the bottom side of $\bar{R}_{\emptyset,{\mathbf{q}}}(x)$, we move left until we are $x$ units away from $\bar{O}$). Again, we define $\bar{R}_{\emptyset,\emptyset}(x)=\emptyset$ for all $x$.
@@ 4.3. Two families of polynomials and their connection to the dented regions
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \times\prod_{i=1}^{n}\frac{(x+i)_{m}}{(x+i+1/2)_{m}}\prod_{i=1}^{m}(2x+n+i+2)_{n+i-1}$
@@
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \times\prod_{i=1}^{m}\frac{(x+i)_{n}}{(x+i+1/2)_{n}}\prod_{i=1}^{n}(2x+m+i+1)_{m+i}.$
@@
We are now ready to define our polynomials $P_{\mathbf{l},\mathbf{q}}(x)$ and $\bar{P}_{\mathbf{l},\mathbf{q}}(x)$. For lists of strictly increasing positive integers ${\mathbf{l}}=(l_{1},\ldots,l_{m})$ and ${\mathbf{q}}=(q_{1},\ldots,q_{n})$, $P_{\mathbf{l},\mathbf{q}}(x)$ and $\bar{P}_{\mathbf{l},\mathbf{q}}(x)$ are given by  [^17] These are equations (5.1) and (5.2) in [5, Part B].181818 For $m=0$, one needs to take $l_{m}=0$ in equations (4.1) and (4.2); also, when $m$ or $n$ is zero at the upper limit of an outside product, that double product is defined to be 1.
@@
Let $H$ be one of the regions $H_{\mathbf{l}}(a,b,k)$, $H_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)$ or $\bar{H}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)$. Then we have191919 Here $H^{+}$ and $H^{-}$ are the regions obtained from the symmetric region $H$ by the procedure described before the statement of Theorem 2.1.
@@
and all of the regions $H^{+}$ and $H^{-}$ are $R$- or $\bar{R}$-regions, as follows202020 Given two regions $R$ and $Q$, $[R,Q]$ is the (ordered) list consisting of these regions.:
((a)). For $k$ even, we have
$H_{\mathbf{l}}(a,b,k)^{+},H_{\mathbf{l}}(a,b,k)^{-}]=\begin{cases}[\bar{R}_{{\mathbf{l}}-1,\emptyset}(a/2),R_{\emptyset,{\mathbf{l}}}((a+k-2)/2)],&\text{$a$ even, $l_{1}=1$,}\\ [{R}_{{\mathbf{l}}-1,\emptyset}(a/2),R_{\emptyset,{\mathbf{l}}}((a+k-2)/2)],&\text{$a$ even, $l_{1}>1$,}\\ [\bar{R}_{{\mathbf{l}},\emptyset}((a-1)/2),R_{\emptyset,{\mathbf{l}}^{(m)}}((a+k-1)/2)],&\text{$a$ odd,}\end{cases}$ (4.6)
where ${\mathbf{l}}-1$ denotes the list obtained from $\mathbf{l}$} by decrementing each of its elements by one unit ((discarding the first element if $l_{1}=1$)).
((b)). For $k$ odd, we have
$H_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)^{+},H_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)^{-}]=\begin{dcases}[{R}_{{\mathbf{q}},{\mathbf{l}}^{(m)}}(a/2),\bar{R}_{{\mathbf{l}},{\mathbf{q}}}((a+k-1)/2)],&\text{$a$ even,}\\ [{R}_{{\mathbf{q}},{\mathbf{l}}}((a-1)/2),\bar{R}_{{\mathbf{l}},{\mathbf{q}}^{(n)}}((a+k)/2)],&\text{$a$ odd,}\end{dcases}$ (4.7)
@@
[…]
@@
Let $H$ be one of the regions $H^{\prime}_{\mathbf{l}}(a,b,k)$, $H^{\prime}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)$ or $\bar{H}^{\prime}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)$, and let $\ell^{\prime}$ be the axis with respect to which the holes in $H$ are symmetric. Then we have212121 Here ${H}^{+}$ and ${H}^{-}$ are the “half-regions” obtained from the nearly symmetric hexagon with holes $H$ by the procedure described before the statement of Theorem 2.2 (see the middle picture in Figure 25 for an example).
@@
((a)). For $k$ even, we have
$H^{\prime}_{\mathbf{l}}(a,b,k)^{+},H^{\prime}_{\mathbf{l}}(a,b,k)^{-}]=\begin{cases}[\bar{R}_{{\mathbf{l}}-1,\emptyset}((a-1)/2),R_{\emptyset,{\mathbf{l}}}((a+k-1)/2)],&\text{$a$ odd, $l_{1}=1$,}\\ [R_{{\mathbf{l}}-1,\emptyset}((a-1)/2),R_{\emptyset,{\mathbf{l}}}((a+k-1)/2)],&\text{$a$ odd, $l_{1}>1$,}\\ [\bar{R}_{{\mathbf{l}},\emptyset}((a-2)/2),R_{\emptyset,{\mathbf{l}}^{(m)}}((a+k)/2)],&\text{$a$ even.}\end{cases}$ (4.10)
((b)). For $k$ odd, we have
$H^{\prime}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)^{+},H^{\prime}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)^{-}]=\begin{dcases}[{R}_{{\mathbf{q}},{\mathbf{l}}^{(m)}}((a-1)/2),\bar{R}_{{\mathbf{l}},{\mathbf{q}}}((a+k)/2)],&\text{$a$ odd,}\\ [{R}_{{\mathbf{q}},{\mathbf{l}}}((a-2)/2),\bar{R}_{{\mathbf{l}},{\mathbf{q}}^{(n)}}((a+k+1)/2)],&\text{$a$ even,}\end{dcases}$ (4.11)
@@
[…]
@@
Let $H$ be one of the regions $H^{\prime}_{\mathbf{l}}(a,b,k)$, $H^{\prime}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)$ or $\bar{H}^{\prime}_{{\mathbf{l}},{\mathbf{q}}}(a,b,k)$, and let $\ell^{\prime}$ be the axis with respect to which the holes in $H$ are symmetric. Then we have222222 Here $\widehat{H}^{+}$ and $\widehat{H}^{-}$ are the “half-regions” obtained from the nearly symmetric hexagon with holes $H$ by the procedure described before the statement of Theorem 2.2 (see the middle picture in Figure 26 for an example).